# MeRF: Motivation-enhanced Reinforcement Finetuning for Large Reasoning Models

### v1

@@ 4.2 Main Results
To demonstrate the effectiveness of MeRF in the reinforcement finetuning process, we compare the performance of MeRF with the RLVR baseline on the K&K logic puzzle dataset. The results in Figure 1 show that MeRF achieves a significant improvement over the RLVR baseline in the validation accuracy within the training process, revealing the remarkable effectiveness of the motivation-enhanced learning for reinforcement finetuning, with only injecting the in-context motivation. In addition, we also observe that the response length of MeRF is comparable with the RLVR models, indicating our method of more efficient reasoning capability.
@@
The results in Table 1 present the validation accuracy in the evaluation of different difficulty levels, comparing the performance of MeRF with the RLVR baseline, startpoint model, and other well-known models. MeRF achieves a significant improvement on the logic reasoning tasks from the startpoint Qwen2.5-7B-Instruct, with only hundreds of training steps, outperforming the RLVR baseline and even some commercial models in all difficulty levels including OOD scenarios. The results of the other baseline models suggest that K&K logic puzzles are challenging tasks for LLMs, and unseen in the training of most models, which further validates the fitness of K&K logic puzzles for analyzing the reasoning capabilities of RLVR models, and the effectiveness of MeRF in the reinforcement finetuning process.
5 Analysis
In this section, we further analyze the effectiveness of MeRF and the impact of motivation on the performance. We conduct the experiments with Qwen2.5-3B-Instruct and Qwen2.5-7B-Instruct as the startpoint models for MeRF and RLVR baseline. We additionaly conduct the experiments of SFT on both models to analyze the performance of SFT with motivation. The results are shown in Table 2.
Table 2: Performance comparison across models on tasks with varying difficulty by number of people.
Model Difficulty by Number of People Avg.
3 4 5 6 7 Avg. 2 (OOD) 8 (OOD)
Qwen2.5-3B-Instruct 0.02 0.00 0.01 0.00 0.00 0.01 0.17 0.00 0.03
   +SFT (w/o motivation) 0.64 0.50 0.25 0.21 0.15 0.35 0.81 0.10 0.38
   +SFT (w/ motivation) 0.28 0.27 0.22 0.18 0.10 0.21 0.34 0.10 0.21
   +RLVR (Baseline) 0.43 0.37 0.20 0.13 0.08 0.24 0.54 0.09 0.26
   +MeRF (Ours) 0.21 0.21 0.19 0.09 0.08 0.16 0.47 0.06 0.19
Qwen2.5-7B-Instruct 0.24 0.10 0.06 0.04 0.04 0.10 0.43 0.00 0.13
   +SFT (w/o motivation) 0.38 0.22 0.14 0.07 0.10 0.18 0.55 0.06 0.22
   +SFT (w/ motivation) 0.54 0.39 0.27 0.13 0.12 0.29 0.60 0.10 0.31
   +RLVR (Baseline) 0.68 0.67 0.57 0.43 0.22 0.51 0.71 0.28 0.51
   +MeRF (Ours) 0.78 0.73 0.68 0.62 0.42 0.65 0.76 0.39 0.63
Figure 3: (1) Left: MeRF underperforms RLVR baseline for Qwen2.5-3B-Instruct. (2) Right: The validation accuracy gap of MeRF is mainly caused by the gap between in-context motivation training and non-motivation validation.
Q1: How does MeRF perform in different models?
As shown in Table 2, the results show that MeRF, on 3B model, underperforms RLVR baseline, while on 7B model, MeRF achieves a significant improvement over RLVR baseline. The validation accuracy of the training curve in Figure 3 (Left) further demonstrates the performance gap between RLVR baseline and MeRF, validating the failure of in-context motivation training for 3B model. In Figure 3 (Right), we observe that when 3B-MeRF is validated with in-context motivation, it achieves comparable performance with 3B-RLVR, and the performance drop is mainly caused by the gap between in-context motivation training and non-motivation validation, shwocasing the 3B model is poor in generalizing to non-motivation validation when trained with in-context motivation. Moreover, we compare it with 7B model in Figure 4, and there is negligible performance drop for 7B model with non-motivation validation, indicating the generalizing capability of model is essential for MeRF to be effective, suggestting that larger models provide the necessary capacity for the method to succeed with MeRF, which is consistent with the findings on the success of RLVR (Yue et al., 2025).
Figure 4: There is negligible performance drop of 7B-MeRF model in non-motivation validation, while 3B-MeRF model fails to generalize to non-motivation validation.
Q2: Can SFT benefit from motivated-enhanced training?
In Table 2, we observe that for 3B model, SFT (w/o motivation) achieves the best performance compared to SFT (w/ motivation), RLVR and MeRF, suggesting that high-quality supervised training with aligned training and validation settings is a more effective strategy for enhancing the performance of small models, which is consistent with the findings of (Guo et al., 2025). The performance drop of the validation gap demonstrated in Figure 3 (Right) also indicates that the model is poor in generalizing to non-motivation validation when SFT with in-context motivation. For 7B model, SFT (w/ motivation) achieves better performance than SFT (w/o motivation), demonstrating that the motivation is beneficial for the performance of SFT, and the model is capable of generalizing to non-motivation validation. However, MeRF achieves the best performance with RLVR baseline outperforming SFT, demonstrating the superiority of reinforcement finetuning for LLMs reasoning (Chu et al., 2025).
@@ 5 Analysis
Q3: Does the performance improvement come from the in-context inference?
@@
We conduct experiments to investigate where the performance improvement mainly comes from. As results shown in Figure 2 (Left), we observe that for both RLVR and MeRF model, they achieve a slightly better performance in validation with motivation compared to non-motivation validation. However, comparing to the performance improvement of MeRF over RLVR in both non-motivation and motivation validation, we find that the performance improvement of MeRF is mainly from the motivation-enhanced reinforcement finetuning process, demonstrating the effectiveness of the in-context motivation in the training process. The results in Figure 4 also show that the performance of 7B-MeRF does not benefit much from the in-context motivation validation, indicating that the performance improvement of MeRF is not mainly from the in-context inference.
Q4: How do the different motivations affect the performance?
@@
Figure 5: (1) Left: The performance of MeRF with different motivations. The motivation with ground-truth reward function achieves the best performance. Adverse motivation misleads the model to provide the wrong answer, while the model is capable of adapting to the adverse motivation in the process of reinforcement finetuning. (2) Right: Examples of different motivations. Suboptimal motivation only describes the correctness score, while adverse motivation provides the full description of the reward function with opposite scores.
The results demonstrate that the motivation with ground-truth reward function achieves the best performance, and the suboptimal motivation performs better than the RLVR baseline, with additional Correctness Score description included in the motivation. Adverse motivation provides the full description of the reward function as ground-truth motivation, while all the score is assigned to the opposite, which misleads the model to provide the wrong answer. The performance drop of the adverse motivation is mainly caused by the contradiction between the in-context motivation and the reward function of the optimization process. However, after several rounds of unstable learning dynamics, the model adapted to discount the motivation signal, treating the score as uninformative, while the full description of the reward function is still beneficial for the model to outperform the RLVR baseline and Motivation (Suboptimal). The results suggest that MeRF benefits from a better alignment between the motivation and the underlying reward function, resulting in better performance, while the model is capable of adapting to the adverse motivation in the process of reinforcement finetuning.
@@ 6 Conclusion
In this paper, we propose a novel method MeRF, leveraging the in-context abilities of LLMs to enhance reinforcment finetuning for reasoning. By injecting the in-context motivation into the training process, MeRF enables the model to be aware of the objective of the task, which aligns the generation with the transcendent optimization objective. Empirical results show that MeRF significantly outperforms RLVR baseline with superior reasoning capabilities. To further understanding the effectiveness of MeRF, we conduct a series of experiments and analysis, revealing the connection between the in-context motivation and the reinforcement learning process optimized by the reward function, providing insights into the potential of motivation-enhanced learning for LLM reasoning and self-evolving LLMs.
@@ Appendix A Appendix
A.1 Broader Impacts
This paper proposes a novel method MeRF for LLM reasoning, leveraging the in-context abilities of LLMs to enhance reinforcement finetuning for reasoning. The proposed method has the potential to improve the performance of LLMs in various reasoning tasks, which can have positive impacts on applications such as natural language understanding, question answering, and logical reasoning. However, it is important to consider how to ensure reliability and safety of the emerged reasoning capabilities of LLMs.
@@ A.2 Related Work
Reinforcement Learning for LLMs. Reinforcement learning (RL) has become a powerful paradigm for fine-tuning LLMs, first demonstrated by reinforcement learning from human feedback (RLHF) to align models with human-preferred responses [Ouyang et al., 2022, Rafailov et al., 2023]. More recently, inspired by the success of DeepSeek-R1 [Guo et al., 2025], Reinforcement Learning with Verified Reward (RLVR) [Lambert et al., 2024] has been proposed to directly enhance reasoning performance by rewarding verifiable success criteria instead of subjective human judgments, which is more similar to the traditional RL paradigm. RLVR has shown promising results in improving the performance of LLMs on various reasoning tasks, including logical reasoning [Xie et al., 2025], coding [Le et al., 2022, Liu et al., 2023], and math problems [Zeng et al., 2025], with simple, designed reward functions. However, these approaches demonstrate the potential of the conventional RL paradigm for LLMs, which may not fully leverage the in-context learning abilities of LLMs, the key to the success of LLMs in previous works [Wei et al., 2022, Brown et al., 2020]. In this work, we propose a novel method MeRF, leveraging the in-context abilities of LLMs to enhance reinforcement finetuning for reasoning.
@@
In-context Learning. In-context learning (ICL) refers to the ability of large language models to learn a task purely from examples and instructions provided in the prompt, without any gradient-based updates to the parameters [Brown et al., 2020]. This ability has been shown to be effective in various tasks, including few-shot learning, zero-shot learning, and even one-shot learning [Wei et al., 2022]. ICL is a key feature of LLMs, enabling them to generalize from a few examples and adapt to new tasks quickly. The success of ICL has led to the development of various prompting strategies, such as few-shot prompting, chain-of-thought prompting, and self-consistency prompting [Wang et al., 2022]. In this work, we investigate the potential of ICL for reinforcement finetuning of LLMs, and propose a novel method MeRF to leverage the in-context abilities of LLMs to enhance reinforcement finetuning for reasoning, injecting the in-context motivation into the training process.
A.3 More Experimental Results
How does MeRF perform in larger models?
To further analyze the performance of MeRF in larger models, we conduct experiments with Qwen2.5-14B-Instruct as the startpoint model. The results are shown in Figure 6, where we observe that MeRF achieves a significant improvement over RLVR baseline in the validation accuracy within the training process, revealing the remarkable effectiveness of the motivation-enhanced learning for reinforcement finetuning, with only injecting the in-context motivation. The results further validate our findings in the Section 5 that larger models provide the necessary capacity for the method to succeed with MeRF, and demonstrate the effectiveness of MeRF in the reinforcement finetuning process.
Case Study.

### v2

@@ 3.2 Main Results
To demonstrate the effectiveness of MeRF in the reinforcement finetuning process, we compare the performance of MeRF with the RLVR baseline on the K&K Logic Puzzle. The results in Figure 1 and Figure 3 show that MeRF consistently achieves a significant improvement over the RLVR baseline in the validation accuracy during the training process, across different model sizes and model families, revealing the remarkable effectiveness of the motivation-enhanced reinforcement finetuning, with simply injecting the in-context motivation.
@@
The results in Table 1 present the validation accuracy in the evaluation of different difficulty levels, comparing the performance of MeRF with the RLVR baseline, startpoint model, and other well-known models. MeRF achieves a significant improvement on the logic reasoning tasks from the startpoint Qwen2.5-7B-Instruct, with only hundreds of training steps, outperforming the RLVR baseline and even some commercial models in all difficulty levels including OOD scenarios. The results of the other baseline models suggest that K&K logic puzzles are challenging tasks for LLMs, and unseen in the training of most models, which further proves the fitness of K&K logic puzzles for analyzing the reasoning capabilities of RLVR models in our experiments, and the effectiveness of MeRF in the reinforcement finetuning process.
Table 2 shows the performance comparison between MeRF and the RLVR baseline on MATH benchmarks, with Qwen2.5-7B-Base model as the startpoint. We report the pass@$k$ performance for $k\in\{1,2,4,8\}$ on AIME24&25, AMC23, and MATH500 datasets, where MeRF achieves consistent improvements over the RLVR baseline in most metrics across all datasets, with an average gain of 3.40%, 1.13%, 3.50%, and 4.50% in pass@1, pass@2, pass@4, and pass@8, respectively. The results demonstrate the effectiveness of MeRF in enhancing the reasoning capabilities of LLMs in the reinforcement finetuning process on mathematical tasks.
Figure 3: Validation Accuracy of MeRF and RLVR baseline on two DeepSeek-R1-Distill models of different sizes and from different model families on K&K. MeRF consistently outperforms the RLVR baseline, demonstrating the effectiveness and generalization of MeRF.
Table 2: Comparison between RLVR baseline and MeRF on math reasoning datasets. We report pass@$k$ for $k\!\in\!\{1,2,4,8\}$ on AIME24&25, AMC23, and MATH500 datasets. Best results are in bold; Average block highlights overall performance and shows absolute gains over Baseline in parentheses. MeRF consistently outperforms the RLVR baseline across these datasets and metrics.
Method AIME24 AIME25 AMC23
pass@1 pass@2 pass@4 pass@8 pass@1 pass@2 pass@4 pass@8 pass@1 pass@2 pass@4 pass@8
RLVR (Baseline) 16.7 20.0 20.0 26.7 6.7 16.7 16.7 20.0 47.5 57.5 70.0 72.5
MeRF (Ours) 20.0 20.0 26.7 30.0 6.7 10.0 16.7 26.7 55.0 67.5 72.5 77.5
[…]
Figure 4: Entropy of models during the training process. MeRF maintains a higher entropy than the RLVR baseline, indicating that MeRF encourages more exploration by the in-context motivation during the training process, which contributes to its improved performance.
4 Analysis on the mechanism behind MeRF
In this section, we further analyze the effectiveness of MeRF by answering the following questions:
Figure 5: Pass@$k$ performance of MeRF and RLVR baseline during the training process (from 0 to 280 steps) on K&K Logic Puzzle. We compare the pass@1, pass@2, pass@4, and pass@8 performance at each step, where MeRF consistently outperforms the RLVR baseline in all metrics. More importantly, MeRF demonstrates a significant training efficiency over RLVR baseline, for example, achieving better pass@4 and pass@8 performance at step 140 than the final RLVR model (at step 280), while RLVR’s performance of pass@4 and pass@8 hardly improves after step 140.
Figure 6: Validation accuracy, Entropy, and Pass@4 on CountDown dataset. MeRF outperforms the RLVR baseline with better accuracy and pass@4 performance, while maintaining a higher entropy during the training process.
@@ 4 Analysis on the mechanism behind MeRF
Q1: Does the performance improvement of MeRF come from the in-context inference?
@@
We conduct experiments to investigate where the performance improvement mainly comes from. As shown in Figure 2 (Right), we observe that for both RLVR and MeRF model, motivation validation leads to slightly better performance (4% and 2%) than non-motivation validation. However, compared to the performance improvement of MeRF over RLVR in both non-motivation validation and motivation validation (27% and 25%), the performance improvement of MeRF is mainly from the motivation-enhanced reinforcement finetuning process, demonstrating the effectiveness of the in-context motivation in the training process. The results in Figure 8 also show that the performance of both models does not benefit much from the in-context motivation validation, indicating that the performance improvement of MeRF is not from the in-context inference.
Q2: If the performance improvement is not from in-context inference, how does the in-context motivation help to enhance the reinforcement finetuning process?
Figure 7: Comparison of pass@8 and pass@1 performance of MeRF and RLVR baseline on MATH500 dataset during the training process. MeRF outperforms the RLVR baseline consistently in both pass@8 and pass@1 metrics, while RLVR pass@8 performance hardly improves after step 80, demonstrating the effectiveness of MeRF in improving the math reasoning capabilities of LLMs.
To further understand how the in-context motivation helps the reinforcement finetuning process, we analyze the pass@$k$ performance and entropy of the models during the training process. The results in Figure 5 show that MeRF consistently outperforms the RLVR baseline in all pass@$k$ metrics, demonstrating the effectiveness of MeRF in improving the reasoning capabilities of LLMs. More importantly, while RLVR’s performance of pass@4 and pass@8 hardly improves after step 140, MeRF achieves better pass@4 and pass@8 performance at step 140 than the final RLVR Model (at step 280). The results in Figure 7 also show that MeRF outperforms the RLVR baseline consistently in both pass@8 and pass@1 metrics on the MATH500 dataset, while RLVR pass@8 performance hardly improves after step 80. Pass@$k$ metrics serve as an indicator of the model’s ability to explore diverse reasoning paths and reach a correct answer (Zeng et al., 2025; Shao et al., 2024). Better pass@$k$ performance and continuous pass@$k$ improvement during the training process suggest the model is more likely to reach a positive reward for optimization and performance improvement. As 8 is the rollout group size of GRPO in our experiments, the growing improvement of pass@8 performance of MeRF over the RLVR baseline indicates that training process progressively amplifies the initial pass@8 improvement with better exploration ability, initially benefiting from the in-context motivation, which also explains why the motivation validation only leads to slightly better performance than non-motivation validation but significant performance improvement of MeRF over RLVR baseline in Figure 2 (Right). The results in Figure 4 show that MeRF maintains a higher entropy than the RLVR baseline during the training process, indicating that MeRF encourages more exploration by the in-context motivation, avoiding the entropy collapse issue of RL for LLMs (Cui et al., 2025), which contributes to its improved performance. The results in Figure 6 also show that MeRF outperforms the RLVR baseline with better accuracy and pass@4 performance on the CountDown dataset, while maintaining a higher entropy during the training process. The results suggest that MeRF benefits from the exploration ability encouraged by the in-context motivation, leading to better performance in the reinforcement finetuning process.
Takeaway: The performance improvement of MeRF is not mainly from the in-context inference, but mainly from the training process. Training process progressively amplifies the initial pass@k improvement with better exploration ability, initially benefiting from the in-context motivation.
Q3: Does the training and validation gap (validation without motivation) affect the performance?
Figure 8: Validation accuray w/ motivation and w/o motivation in the prompt. Both models achieve comparable performance between two validation settings, indicating the negligible impact of the training and validation gap.
As MeRF includes the in-context motivation in the training process, while the validation is conducted without motivation, there exists a training and validation gap, which may affect the performance of MeRF. To investigate the impact of the training and validation gap on the performance, we conduct experiments to validate the models with and without motivation.
The results are shown in Figure 8, where we compare the performance of Qwen-2.5-7B-Instruct and DeepSeek-R1-Distill-Qwen-1.5B (Guo et al., 2025) trained with MeRF in the two validation settings. We find that both models achieve comparable performance between the two validation settings, indicating the negligible impact of the training and validation gap on the performance of these models. The results suggest that the model is capable of generalizing to non-motivation validation when trained with in-context motivation, which is essential for MeRF to be effective.
Q4: How do the different motivations (suboptimal, adverse) affect the performance?
@@
Figure 9: (Left) The performance of MeRF with different motivations. The motivation with ground-truth reward function achieves the best performance. Adverse motivation misleads the model to provide the wrong answer, while the model is capable of adapting to the adverse motivation in the process of reinforcement finetuning. (Right) Examples of different motivations. Suboptimal motivation only describes the correctness score, while adverse motivation provides the full description of the reward function with opposite scores.
The results demonstrate that the motivation with ground-truth reward function achieves the best performance, and the suboptimal motivation performs better than the RLVR baseline, with an additional correctness score description included in the motivation. Adverse motivation provides the full description of the reward function same to ground-truth motivation, while all the score is assigned to the opposite, which misleads the model to provide the wrong answer. The performance drop of the adverse motivation is mainly caused by the contradiction between the in-context motivation and the reward function of the optimization process. However, after several rounds of unstable learning dynamics, the model adapted to discount the motivation signal, either treating the score as uninformative, or understanding the deliberately adverse motivation and opposite scores, while the full description of the reward function is still beneficial for the model to outperform the RLVR baseline and Motivation (Suboptimal). The results suggest that MeRF benefits from a better consistency between the motivation and the underlying reward function, resulting in better performance, while the model is capable of adapting to the adverse motivation in the process of reinforcement finetuning.
Takeaway: The training and validation gap caused by in-context motivation has a negligible impact on the performance of models with strong generalization capabilities. MeRF benefits from a better consistency between the motivation and the underlying reward function, while the model is capable of adapting to the adverse motivation in the process of reinforcement finetuning.
@@ 5 Conclusion
In this paper, we propose MeRF, leveraging the in-context learning abilities of LLMs for more efficient reinforcement finetuning with in-context motivation. By injecting the in-context motivation into the training process, MeRF enables the model to be aware of the objective of the task, aligning the generation with the transcendent optimization objective, and therefore, leading to substantial performance improvement in reasoning benchmarks. To further understand the effectiveness of MeRF, we conduct comprehensive experiments and analysis, revealing the mechanism behind MeRF, demonstrating the powerful capability of LLMs in adapting to adverse motivation and the potential for more powerful large reasoning models with motivation-enhanced reinforcement finetuning.
@@
Ethics Statements
This work adheres to the ICLR Code of Ethics. Our research does not involve human subjects, personally identifiable information, or sensitive personal data. All datasets used are publicly available and commonly adopted in prior work; we followed their intended licensing and usage guidelines. We have carefully considered potential risks of misuse, including issues of fairness, bias, and safety, and we believe our contributions do not raise immediate ethical concerns.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work. The main paper provides a complete description of the proposed method and experimental setup. The appendix includes additional details on the training procedures, and hyperparameter settings. We will release the code, and detailed instructions for reproducing our experiments upon acceptance.
@@ Appendix A Appendix
A.1 Use of LLMs
We used Large Language Models (LLMs) as a general-purpose writing and editing tool to improve grammar, clarity, and readability of the manuscript. LLMs were not used for research ideation, experiment design, data analysis, or derivation of results. All technical contributions, experimental designs, and scientific claims were developed by the authors.
@@ A.2 Related Work
Reinforcement Learning for LLMs. Reinforcement learning (RL) has become a powerful paradigm for fine-tuning LLMs, first demonstrated by reinforcement learning from human feedback (RLHF) to align models with human-preferred responses (Ouyang et al., 2022; Rafailov et al., 2023). More recently, inspired by the success of DeepSeek-R1 (Guo et al., 2025), Reinforcement Learning with Verified Reward (RLVR) (Lambert et al., 2024) has been proposed to directly enhance reasoning performance by rewarding verifiable success criteria instead of subjective human judgments, which is more similar to the traditional RL paradigm. RLVR has shown promising results in improving the performance of LLMs on various reasoning tasks, including logical reasoning (Xie et al., 2025), coding (Le et al., 2022; Liu et al., 2023), and math problems (Zeng et al., 2025), with simple, designed reward functions. However, these approaches demonstrate the potential of the conventional RL paradigm for LLMs, which may not fully leverage the in-context learning abilities of LLMs, the key to the success of LLMs in previous works (Wei et al., 2022; Brown et al., 2020). In this work, we propose a novel method MeRF, leveraging the in-context abilities of LLMs to enhance reinforcement finetuning for reasoning.
@@
In-context Learning. In-context learning (ICL) refers to the ability of large language models to learn a task purely from examples and instructions provided in the prompt, without any gradient-based updates to the parameters (Brown et al., 2020). This ability has been shown to be effective in various tasks, including few-shot learning, zero-shot learning, and even one-shot learning (Wei et al., 2022). ICL is a key feature of LLMs, enabling them to generalize from a few examples and adapt to new tasks quickly. The success of ICL has led to the development of various prompting strategies, such as few-shot prompting, chain-of-thought prompting, and self-consistency prompting (Wang et al., 2022). In this work, we investigate the potential of ICL for reinforcement finetuning of LLMs, and propose a novel method MeRF to leverage the in-context abilities of LLMs to enhance reinforcement finetuning for reasoning, injecting the in-context motivation into the training process.
A.3 Preliminary
Reinforcement Learning with Verifiable Rewards (RLVR) (Lambert et al., 2024) is a reinforcement learning paradigm for training language models on tasks with verifiable outcomes, such as math problems or logic puzzles. The key idea is to use a reward function that can be automatically verified, allowing the model to learn from the ultimate correctness of its outputs. Recent works (Zeng et al., 2025; Yu et al., 2025) have shown that RLVR can significantly improve the reasoning capabilities of LLMs, with reasoning patterns emerging from the optimization for verifiable rewards.
Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is a reinforcement learning algorithm designed to optimize policies by leveraging group-wise relative preference information. As a variant of Proximal Policy Optimization (PPO) (Schulman et al., 2017), GRPO foregoes the need for critic models and instead focuses on learning from relative comparisons of actions within groups, significantly enhancing training efficiency for reinforcement finetuning of LLMs. For each question $x$, GRPO samples a group of $G$ outputs $\{y_{i}\}_{i=1}^{G}$ from the policy $\pi_{\theta_{\text{old}}}(\cdot|x)$, and computes the advantage $A_{i}$ for each output $y_{i}$ based on the outcome reward $r_{i}$, where $\pi_{\text{ref}}$ is a reference model. The GRPO objective is defined as follows:
$\mathcal{L}_{\text{GRPO}}(\theta)$ $=\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\left[\frac{1}{G}\sum_{i=1}^{G}\min\left(\rho_{i}A_{i},\text{clip}(\rho_{i},1-\varepsilon,1+\varepsilon)A_{i}\right)\right]$
$\quad-\beta\mathbb{D}_{\text{KL}}\left(\pi_{\theta}||\pi_{\text{ref}}\right),$ (1)
where
$\rho_{i}=\frac{\pi_{\theta}(y_{i}|x)}{\pi_{\theta_{\text{old}}}(y_{i}|x)}$ (2)
is the importance ratio and the advantage is computed as:
$A_{i}=\frac{r_{i}-\text{mean}(\{r_{1},r_{2},\cdots,r_{G}\})}{\text{std}(\{r_{1},r_{2},\cdots,r_{G}\})}$ (3)
This normalizes the outcome rewards of the group of outputs and sets the advantage for all the tokens in the output $\{y_{i}\}_{i=1}^{G}$. This formulation enables GRPO to learn from relative preferences within each group without the need for a critic model, making it efficient and suitable for the implementation of RLVR to LLMs.
The Knights and Knaves (K&K) (Xie et al., 2024) logic puzzle dataset is a widely used (Xie et al., 2025) benchmark for reinforcement finetuning for LLMs reasoning, which provides a well-structured difficulty level of the logic tasks and allows accurate and easy reward verification for RLVR. The controllable difficulty levels are achieved by varying the number of people in the logic task, the more people in the logic puzzle requiring LLMs’ more complex reasoning, and more steps to solve the task. The K&K dataset contains 7 different difficulty levels of logic puzzles, with 2 people as the easiest level and 8 people as the most difficult level. Here is an example of the K&K dataset with 3 people:
CountDown (Pan et al., 2025) is a challenging numerical reasoning dataset that requires models to perform arithmetic operations and logical reasoning to arrive at the correct answer. The dataset consists of problems that involve a series of numbers and a target number, where the goal is to use the given numbers and basic arithmetic operations (addition, subtraction, multiplication, and division) to reach the target number. In our experiments, each problem provides a set of 3 or 4 numbers and a target number, and the model must determine a sequence of operations that will result in the target number. The problems in the CountDown dataset vary in difficulty, with some requiring simple calculations while others necessitate more complex reasoning and multiple steps to solve. The dataset is designed to test the model’s ability to understand numerical relationships, perform calculations accurately, and apply logical reasoning to achieve the desired outcome. The CountDown dataset is widely used as a benchmark for evaluating the numerical reasoning capabilities of language models.
An Example of CountDown dataset <|im_start|>system You are a helpful assistant. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think><answer> answer here </answer>. Now the user asks you to solve a math reasoning problem. After thinking, when you finally reach a conclusion, clearly state the equation within <answer> </answer> tags. i.e., <answer> (1 + 2) / 3 </answer>. Now, the user will give you the math reasoning problem to solve. Think carefully, follow the structure. <|im_end|> <|im_start|>user Using the numbers [2, 2, 2], create an equation that equals 8. You can use basic arithmetic operations (+, -, *, /) and each number can only be used once. <|im_end|> <|im_start|>assistant <think>
Other Implementation Details
We implement our method with VeRL (Sheng et al., 2024) and Logic-RL (Xie et al., 2025) codebase. We conduct the reinforcement finetuning with batch size of 16, a learning rate of 1e-6, 1 and 2 epochs of training. We use the rollout framework of Vllm (Kwon et al., 2023) for sampling the outputs of the model. The group size of GRPO is set to 8, the clip range is set to 0.2, and the KL penalty is set to 0.001.
A.4 Case Study